# “Talking Past Each Other”: Issue Ownership and Microtargeting in Swiss Online Political Ads

### v1

@@ Appendix A DATA
To vote / To elect "wählen", "wähle", "wählen wir" "votare", "vota", "votiamo", "voter", "vote", "votons",
"eleggere", "eleggi", "eleggiamo" "élire", "élis", "élisons"
Election(s) / Vote(s) "Wahl", "Wahlen" "elezione", "elezioni", "élection", "élections",
"votazioni", "voto" "votes", "vote"
Poll / Ballot "Abstimmungen", "Abstimmung" (n/a) (n/a)
Referendum "Referendum" "referendum" "référendum"
Electoral Campaign "Wahlkampf", "Wahlkampagne" "campagna elettorale" "campagne électorale"
Electoral Round "Wahlgang" "turno elettorale" "tour électoral"
Counting / Tally "Auszählung" "scrutinio" "scrutin"
Table A2: Impressions by party and language
@@
Table A3: Number of ads by party and language
@@
Table A4: Expenditure on ads by party and language
@@
Appendix B METHODS
Table A5: Results of the Wilcoxon signed-rank test (p < 0.001).
@@ Appendix B METHODS
Figure A1: Confusion matrices for the Demographic + Topics model.

### v2

@@ 7 Supplementary Material
To vote / To elect “wählen”, “wähle”, “wählen wir” “votare”, “vota”, “votiamo”, “voter”, “vote”, “votons”,
“eleggere”, “eleggi”, “eleggiamo” “élire”, “élis”, “élisons”
Election(s) / Vote(s) “Wahl”, “Wahlen” “elezione”, “elezioni”, “élection”, “élections”,
“votazioni”, “voto” “votes”, “vote”
Poll / Ballot “Abstimmungen”, “Abstimmung” (n/a) (n/a)
Referendum “Referendum” “referendum” “référendum”
Electoral Campaign “Wahlkampf”, “Wahlkampagne” “campagna elettorale” “campagne électorale”
Electoral Round “Wahlgang” “turno elettorale” “tour électoral”
Counting / Tally “Auszählung” “scrutinio” “scrutin”
Table 3: Impressions by party and language
@@
Table 4: Number of ads by party and language
@@
Table 5: Expenditure on ads by party and language
@@
Table 6: Wilcoxon signed-rank test of ad views 30 days pre- and post-referendum ($p<0.001$ for all tests).